A researcher wishes to estimate the proportion of adults who have​ high-speed Internet access. What size sample should be obtained if she wishes the estimate to be within 0.02 with 95​% confidence if ​(a) she uses a previous estimate of 0.42​? ​(b) she does not use any prior​ estimates?
Answer #1

(a)Solution :

Given that,

= 0.42

1 - = 1 - 0.42 = 0.58

margin of error = E = 0.02

At 95% confidence level the z is ,

= 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z0.025 = 1.96

Sample size = n = (Z/2 / E)2 * * (1 - )

= (1.96/ 0.02)2 * 0.42 * 0.58

= 2339.53 = 2340

Sample size = 2340

(b) Solution :

Given that,

= 0.5 (when estimate is not given then 0.5)

1 - = 1 - 0.5 = 0.5

margin of error = E = 0.02

At 95% confidence level the z is ,

= 1 - 95% = 1 - 0.95 = 0.05

/ 2 = 0.05 / 2 = 0.025

Z/2 = Z0.025 = 1.96

Sample size = n = (Z/2 / E)2 * * (1 - )

= (1.96/ 0.02)2 * 0.5 * 0.5

= 2401

Sample size = 2401